OpenStudy (anonymous):

For what x-value(s) does sin(x) = -1?

OpenStudy (ibbutibbu.):

\[x = \frac{ \pi }{ 2 } \pm 2\pi n\]

OpenStudy (anonymous):

Is this the answer, or...?

OpenStudy (ibbutibbu.):

this is the answer.

OpenStudy (ibbutibbu.):

oh wait

OpenStudy (ibbutibbu.):

no

OpenStudy (anonymous):

Okay...

OpenStudy (jdoe0001):

check your Unit Circle

OpenStudy (ibbutibbu.):

\[x = \frac{ 3\pi }{ 2 } \pm 2\pi n\]

OpenStudy (ibbutibbu.):

thats better